# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v1

@@ 3.2.2. Election Forecast Displays
Imagine a study targeting the effect of different ways of visualizing election forecasts. The researchers develop a simple forecasting model to predict vote share (a percentage) by weighting and aggregating poll results from news sources for ten unique upcoming two-party state-level elections. Participants are randomly assigned to view one of three forecast displays showing the predicted vote share with uncertainty of each party’s candidate as a table, an animated graphic, or a static graphic. Assume for simplicity that all three induce informationally equivalent signals to a rational agent [^4] If any of the displays was not informationally equivalent to the others, the optimal decision may vary by display. See Wu et al. (Wu et al., 2023) for analysis of such examples using a rational agent framework..
@@
Upon beginning the experiment, participants are randomly assigned to a display treatment and also assigned a political party preference. They are told they will view election forecasts from a model based on poll results and should imagine they live in the district where the election takes place. Each participant in the study does ten trials corresponding to the ten elections. On each trial, they are asked if they would like to vote for the candidate in the election, elicited using a checkbox listing these options.
@@
Participants are told they will win a given base reward amount plus an additional \$$\$0.25USD$ for each trial where their candidate wins a simulated version of the election that uses the forecast model to simulate how other people will vote (voting carries no cost). The researchers generate outcomes by randomly drawing a bivariate vote share outcome from the predicted distribution under the model and generating the uncertain state $\theta$ (taking the value 0 or 1 for lose or win) by assuming the higher vote share candidate wins. Participants are not given feedback after each trial.
@@
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to the closeness of the race, defined as the difference between the predicted vote share of the first candidate and 0.5. Using statistical significance testing they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types, using posthoc comparisons to declare that the animated graphic leads to better voting decisions because in induces the most sensitivity.
Interpretation of Results. For the researchers’ conclusions to be valid, participants should be given all necessary information to arrive at posterior beliefs informed by $v$ as shown in Equation 3, and to select the utility maximizing action (vote or not) under the scoring rule for the voting task (Equation 2). We first consider whether this is the case.
@@
The action space $A$ ($a=1$ for vote, $a=0$ for not vote) is conveyed to participants, and the task description also gives participants the statespace $\Theta$ ($\theta=1$ for their candidate winning, $\theta=0$ for not). Participants can infer the probability of their candidate winning from the signal showing the model predicted vote share.
@@
What do participants know about the realized outcome of the election that the researchers describe simulating? The researchers described the use of the forecast model to simulate voting behavior, and in implemention, do not account for the participant’s vote in generating the outcome. Consequently, whether a participant votes is irrelevant to the payoff-relevant state under the scoring rule. However, the study design is also somewhat deceptive in that the description they give is ambiguous when it comes to whether the participant’s vote counts: some participants might assume, based on their knowledge of real world elections, that it does.
@@
Because, in the context of the study, voting carries no value but also no cost, we cannot conceive of the four sources of loss. There is no well-defined decision problem, and consequently, it does not make sense to ask whether behavior is more biased or otherwise worse under one display condition.
@@ 3.3. Goals & Incentives in the Lab vs World
Imagine that the researchers who designed the voting task intentionally did not include the vote in the scoring rule, and were intentionally ambiguous about the role of the vote in the data-generating model. Because the utility of voting in a real-world election is typically left implicit and likely to vary between people, the researchers might think that their results would therefore be more representative of real world voting behaviors if the scoring rule was ambiguous.
@@ 4. Applying the Framework to HCAI Studies
We use a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions in the absence of a well-defined decision problem. We randomly sampled 46 of the studies surveyed by Lai et al. (2023) in their overview of empirical human-subjects studies on human-AI decision-making. According to their inclusion criteria, the sample contains only evaluative human-subject studies targeting a decision task in the context of classification or regression published in ACM or ACL conference venues between 2018 and 2021.
@@
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable. As described above, to be applicable to a decision task, the task should allow for defining a ground truth state and induce uncertainty in the payoff-relevant state, at least across participants. Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way. Finally, we identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information to identify the optimal response, at least in principle, such that evaluative conclusions are warranted. When the paper and supplemental materials did not include enough information about the study for us to identify whether the participants had sufficient information (which affected a small handful of cases), we coded it as a negative example.
@@
Two authors participated in the coding. After formulating the coding scheme, the authors divided the sampled studies roughly evenly, each coding independently, then discussing their results to ensure consensus. We provide the list of sampled studies as Appendix.
Results. Of the 46 studies examined, 11 (24%) studied tasks for which ground truth could not be determined, such as subjective appraisals of music or movie recommendations or emotion. One of these studied accuracy, but from a purely descriptive perspective of understanding what accuracy humans versus algorithms could achieve on a task (Lin et al., 2020).
@@
All 35 of the remaining studies drew conclusions about flaws in human decisions, such as by pointing to over- or underreliance on the AI or speculating about why participants failed to get the correct answer. However, participants were only given sufficient information to identify the best decision response across all conditions in 6 of these 35 (17%); in the other 29 cases it was either clear that they did not get sufficient information (24) or only some conditions got sufficient information (5), e.g., through the provision of feedback or model accuracy on held-out data. In the latter case, informational differences between conditions that the researchers compared lead to different a priori expectations of decision quality, yet these differences were not accounted for in comparing participants’ behavior between conditions. This prevents assessing how much of the possible benefit from a treatment participants were able to attain. The calculation of separate performance benchmarks would address such ambiguity (Wu et al., 2023).
@@
Additionally, even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent,and potentially misinterpreting the performance of participants guided by a different scoring rule. In some of these cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model. The most common case was when the researchers gave participants the model’s held-out accuracy on examples different from those participants experienced. When the study design includes feedback, this leads to a contradiction between the instructions and what participants observe.

### v3

@@ 3.3.2. Election Forecast Displays
Imagine a study targeting the effect of different ways of visualizing election forecasts. The researchers develop a forecasting model to predict vote share (a percentage) by weighting and aggregating poll results from news sources for ten unique upcoming two-party state-level elections. Participants are randomly assigned to view one of three forecast displays showing the predicted vote share with uncertainty of each party’s candidate as a table, an animated graphic, or a static graphic. Assume for simplicity that all three induce informationally equivalent signals to a rational agent [^4] If any of the displays was not informationally equivalent to the others, the optimal decision may vary by display. See Wu et al. (Wu et al., 2023) for analysis of such examples using a rational agent framework..
@@
Upon beginning the experiment, participants are randomly assigned to a display treatment and also assigned a political party preference. They are told they will view election forecasts from a model based on poll results and should imagine they live in the district where the election takes place. Each participant in the study does ten trials corresponding to the ten elections. On each trial, they are asked if they would like to vote for the candidate in the election. Voting carries no cost.
@@
Participants are told they will win a given base reward amount plus an additional \$$\$0.25USD$ “for each trial where their candidate wins a simulated version of the election that uses the forecast model to predict how other people will vote.” The researchers generate outcomes by randomly drawing a bivariate vote share outcome from the predicted distribution under the model and generating the uncertain state $\theta$ (taking the value 0 or 1 for lose or win) by assuming the higher vote share candidate wins. Participants are not given feedback after each trial.
@@
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to the closeness of the race, defined as the difference between the predicted vote share of the first candidate and 0.5. The researchers postulate that a better display induces more sensitivity. Using statistical significance testing, they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types, using posthoc comparisons to declare that the animated graphic leads to better voting decisions because it induces the most sensitivity.
@@
Interpretation of Results. To understand why the researchers’ conclusions might be misleading as a prediction of real world use of election displays, consider the decision problem from the perspective of the participants. The action space $A$ ($a=1$ for vote, $a=0$ for not vote) is conveyed to participants, and the task description also gives participants the statespace $\Theta$ ($\theta=1$ for their candidate winning, $\theta=0$ for not). Participants can infer the probability of their candidate winning from the signal showing the model predicted vote share.
@@
What do participants know about the realized outcome of the election that the researchers describe simulating? The researchers described the use of the forecast model to simulate voting behavior, and in implemention, do not account for the participant’s vote in generating the outcome. Consequently, whether a participant votes is irrelevant to the payoff-relevant state under the scoring rule. However, the description provided to participants is ambiguous about whether the participant’s vote counts: some participants might assume, based on their knowledge of real-world elections, that it does. Hence performing well in the experiment requires being able to guess what the researchers have in mind, rather than behaving rationally in the context of the experiment.
@@
Because, in the context of the study, voting carries no value but also no cost, we cannot conceive of the four sources of loss. There is no well-defined decision problem, and consequently, no formal justification for concluding that behavior is more biased or otherwise worse under one display condition.
@@ 3.4. Goals & Incentives in the Lab vs World
Imagine that the researchers who designed the voting task intentionally did not include the vote in the scoring rule, and were intentionally ambiguous about the role of the vote in the data-generating model. Because the utility of voting in a real-world election is typically left implicit and likely to vary between people, the researchers might think that their results would therefore be more representative of real world voting behaviors.
@@ 4. Applying the Framework to HCAI Studies
We use a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions in the absence of a well-defined decision problem. We sampled 46 of the studies surveyed by Lai et al. (2023) in their overview of empirical human-subjects studies on human-AI decision-making. According to their inclusion criteria, the sample contains only evaluative human-subject studies targeting a decision task in the context of classification or regression published in ACM or ACL conference venues between 2018 and 2021.
@@
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable: Does the task given to participants admit a ground truth state and induce uncertainty about the state, at least across participants. Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way. We identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information for a rational agent to best respond to the task. Finally, we noted whether the researchers reported conclusions using a different scoring rule than the one provided to participants, and whether participants were incentivized using a flat scoring rule (i.e., a rule that does not distinguish between responses based on their distance to the ground truth). Note that studies that pay participants a flat reward require that researchers report conclusions using a different scoring rule. Our coding guide is in Appendix 7.2.
In some cases, coding was challenging because the paper did not provide much detail on what exactly was communicated to participants, nor provide the study interface as a supplement. We note such cases in our coding table (Appendix 7.3).
@@
Two authors participated in the coding. After formulating the coding scheme, the authors divided the sampled studies roughly evenly, each coding independently, then discussing all challenging examples to ensure consensus. We provide the list of sampled studies with codes as Appendix.
Results. Of the 46 studies examined, 7 (15%) studied tasks to which the framework could not be applied. These included studies relegated to perceptions (e.g., of fairness, accuracy) and of subjective appraisals of music or movie recommendations or emotion. One study posed questions and drew conclusions about the diagnostic utility of AI recommendations, but elicited only perceived utility (Cai et al., 2019). We coded this example as Applicable.
@@
All 39 of the remaining studies drew conclusions about flaws in human decisions, such as by pointing to over- or underreliance on the AI or speculating about why participants failed to get the correct answer. However, participants were only given sufficient information to identify the best decision response across all conditions in 7 of these 39 (18%), and across some conditions in another 3 of 39 (7.7%). In the latter cases, informational differences between conditions that the researchers compared could lead to different a priori expectations of decision quality, yet these differences were not accounted for in comparing participants’ behavior between conditions [^5] The calculation of separate performance benchmarks would address such ambiguity (Wu et al., 2023).. There was not enough information to determine if participants were presented with a well defined problem in 4 (10%) of the remaining studies.
Only 7 of the 39 (18%) studies to which the framework is applicable used the same scoring rule to analyze the responses as was used to incentivize participants. 25 (64%) did not, one did in at least one experiment, and 6 others (15%) did not report enough detail for us to know. Finally, 14 (36%) of the 39 used a flat rule for incentives, one did in some experiments, 16 (36%) did not, and 7 (18%) did not report enough detail for us to know.
@@
We observed that even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent,and potentially misinterpreting the performance of participants guided by a different scoring rule. In some cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model. The most common scenario was when the researchers gave participants information about the model’s held-out accuracy but the participants experienced instances where the model had a different accuracy. When the study design includes feedback, this leads to a contradiction between the instructions and what participants observe which makes it difficult to draw evaluative claims about behavior.
@@ 6. Limitations & Future Work
Acknowledgements.
We thank Yifan Wu and Ziyang Guo for discussions on these topics, and Dan Weld for feedback on a draft.